# What’s the Weight?

### v1

@@ 3.2 Proposed Estimators
We are interested in drawing inference about the super population of individuals from which we have selected our study participants. Here, the targets of inference are the conditional outcome means, i.e., $\mathbb{E}\left[Y\mid A=a,X\right]$ for $a\in\mathcal{A}$ and their expected difference over $X$, i.e.,
${\rm ACD}=\mathbb{E}_{X}\left[\mathbb{E}\left[Y\mid A=1,X\right]-\mathbb{E}\left[Y\mid A=0,X\right]\right].$ (1)
Our goal is to estimate the individual group specific conditional means, of which ACD is the difference. We make the following assumptions:
@@
    Positivity: $\Pr(A=a\mid X=x)>0\ \forall a\in\mathcal{A}$ and every $x$ such that $f_{X}(x)>0$
@@
    Positivity of Sample Selection: $\Pr(S=1\mid A=a,X=x)>0$ for every $a,x$ such that $f_{A,X}(a,x)>0$
@@
    Weak Selection Exchangeability: $\mathbb{E}[Y\mid A=a,X]=\mathbb{E}[Y\mid A=a,S=1,X]$
We require (i) and (ii) to ensure that our observed data functional, denoted in the next section by $\mu(a)$, is well defined. Condition (iii) allows us to generalize the conditional mean of the outcome from the study sample to the target population. This assumption is less stringent than exchangeability in distribution (i.e, $Y\perp S\mid A,X$), as it does not require the entire conditional distribution of the outcomes to be independent of selection, only their conditional expectations. Depending on the factorization of the joint probability of being sampled $(S=1)$ and belonging to a particular group $(A=a)$, we propose three formulas, which can be estimated using either outcome model–based estimators (i.e., g-computation) or inverse probability weighted estimators:
@@
$\mathbb{E}_{X}\left[\mathbb{E}[Y\mid A=a,S=1,X]\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid X)}\mid S=1\right],$ (2)
@@
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(S=1\mid A=a,X)\Pr(A=a\mid X)}\mid S=1\right],$ (3)
$\mathbb{E}_{X}\left[\frac{AY\Pr(S=1)}{\Pr(A=a\mid S=1,X)\Pr(S=1\mid X)}\mid S=1\right].$ (4)
We will refer to the methods for estimating Equations (2), (3), and (4) as Outcome Modeling and Direct Standardization (OM), Inverse Probability Weighting 1 (IPW1), and Inverse Probability Weighting 2 (IPW2), respectively. For additional details on this derivation, see Appendix A.1.
@@
The assumptions required to estimate the ACD are relatively weaker than those required to identify a causal quantity. Figure 1, Panel (C) gives an example directed acyclic graph (DAG) where a causal target may be warranted, such as the case where $A$ is an environmental exposure, like lead, which can also alter telomere length. We provide a supplementary analysis in which we estimate the causal effect of lead exposure on telomere length for illustration, as a practitioner may also be interested in such causally interpretable estimated differences. If the group variable under study is a treatment or exposure, we show under stronger assumptions that Formulas (2) - (4) can equivalently target the population potential outcome means, denoted by $\mathbb{E}[Y^{a}]\ \forall a\in\mathcal{A}$. Along with the previous assumptions, the following conditions,
  1. (iii*)
    Weak Selection Exchangeability for Potential Outcomes: $\mathbb{E}[Y^{a}\mid A=a,X]=\mathbb{E}[Y^{a}\mid A=a,S=1,X]$
@@
    Stable Unit Treatment Value Assumption (SUTVA): $Y_{i}=Y_{i}^{a}$ for all observations $i=1,\ldots,n$ and for all $A_{i}=a\in\mathcal{A}.$
@@
    Weak Exchangeability: $\mathbb{E}[Y^{a}\mid X]=\mathbb{E}[Y^{a}\mid A=a,X]$,
are sufficient to identify $\mathbb{E}[Y^{a}]$. The additional SUTVA assumption underpins the existence of the counterfactual outcomes, $Y^{a}$. The compact representation in (iv) captures two aspects of SUTVA – no interference, which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and consistency, which states that the potential outcome is well-defined for each treatment level, $a$ (Rubin, 1980). The weak exchangeability assumption allows us to represent the conditional distribution of the unobserved potential outcome using that of the observed potential outcome. As before, we only require a weaker version of this assumption, though strong exchangeability, i.e., $Y^{a}\perp A\mid X$, is often assumed. For additional details and the full identification proof, see Appendix A.2. This is thematically similar to Dahabreh et al. (2020), who study the generalization of causal inferences from individuals in randomized trials. The authors also differentiate between mean exchangeability in $A$ for the randomized population and mean exchangeability in $S$ for the target population in the context of randomized trials.
@@ 3.3 Comparisons to Existing Approaches
The three formulas derived above differ from standard, existing approaches that one might take in practice, which warrants some brief discussion here. Table 1 compares our proposed approaches to such existing methods. As shown, there are two sources of bias if our assumed conceptual diagram holds – bias from covariate imbalance between comparison groups (referred to herein as confounding bias) and unequal selection induced by the design of the complex survey (referred to herein as selection bias), which impacts the generalizability of our within-sample results to a larger, target population. Existing approaches may incorporate either or both balancing and/or generalizability weights to account for these two sources of bias, however if either are foregone, then we expect to obtain biased estimates of our target quantity of interest. For instance, the estimated (unadjusted) effect of race can be either exaggerated or attenuated depending on the direction of the true association between the confounders (e.g., SES), race, and the outcome in simple regression (Approach 1 in Table 1; Williams et al., 2019). Multiple regression (Approach 2) further adjusts for covariates, but unless the outcome model includes the appropriate interaction terms, this approach will suffer from confounding bias as well (Rao, 1971). With inverse probability weighting (Approach 3), the race-covariate relationship is modeled, rather than the covariate–outcome relationship, foregoing the need to specify any race $\times$ covariate interactions (Mansournia and Altman, 2016). However, with all three of these approaches, the target of inference is the within-sample difference in outcomes, which cannot be generalized to a larger population if a probability sample was taken (Shi et al., 2023; Degtiar and Rose, 2023). Survey-weighted regression (Approach 4) accounts for selection bias by accounting for how many individuals in the population that one individual in the sample represents (Lumley and Scott, 2017), allowing for generalizability, but these results will be biased if there is covariate imbalance in the sample (DuGoff et al., 2014; Ridgeway et al., 2015).
@@
Table 1: Comparison of existing approaches for analysis to our proposed approach. Dashes denote that the model/weight is not used.
IPTW = Inverse Probability of Treatment Weighting
$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X)$
@@
A more principled approach would be one that accounts for both confounding and selection bias, as given by Approaches 6 and 7, which utilize a combination of propensity score and survey weighting. These approaches are advocated for by Zanutto (2006); DuGoff et al. (2014) and Ridgeway et al. (2015), respectively, and may be viewed as current analytic standard since they suggest simple survey-weighted generalizations of the usual use of IPTW, either at the first (propensity model) and/or second (outcome model) stage. Though it may seem that either simple extension addresses both sources of bias, there has been little theoretical justification of how weighting one or multiple models within IPTW accomplishes both tasks. This is also true of the ‘naive g-computation’ approach (Approach 8) we study, which weights the predicted conditional expectation for each individual’s outcome and comparison group/race by their survey contribution. Additionally, propensity scores are often seen as in-sample balancing scores (Zanutto, 2006; DuGoff et al., 2014). From this perspective, their generalizability (and corresponding selection bias correction based on survey weights) is often ignored. Survey-weighting both the propensity score and outcome model is reasonable (Ridgeway et al., 2015), however directly combining survey-weighted IPWs and survey weights in a second stage outcome regression may also lead to biased results if additional correction is not made. In our research question, we have an intersecting consideration of both selection and confounding bias; thus, neither source of bias should be disregarded.
@@
From a practical standpoint, one question at hand is: “should we implement survey-weighted propensity score estimation in IPTW or not”. Our analytic results show that for estimation of the ACD, the propensity score does not have to be survey-weighted. We see this in Equations (3) and (4), which differ with respect to the factorization of the joint distribution into propensity and survey weight models. If the propensity score is not survey-weighted, based on Equation (4), we need $\Pr(S=1\mid X)$, i.e., one may need to marginalize $\Pr(S=1\mid A,X)$ over the group membership variable, $A$. If survey weights are included in the propensity score estimation, then we must know $\Pr(S=1\mid A,X)$, i.e., the probability of selection conditional on each level of the group membership variable and the covariates.
@@
Moreover, our proposed approaches carry additional differences amongst themselves, and prove to be useful in different settings. Formula (2), outcome modeling and direct standardization (OM), requires individual-level data on $X$, or knowledge of $f_{X}(x)$, as an outcome regression, $g_{a}(X)$, is fit in practice. Formulas (3) and (4), IPW1 and IPW2, do not require estimation of an outcome model and may be the best options depending on how much information is available in practice. We view IPW1 and IPW2 as simple corrections to what one might do in practice, namely the traditional inverse probability of treatment weighted (IPTW) estimator, and thus may be a better starting point for an analysis of this type as they additionally account for selection bias. Further, IPW1 and IPW2 do not assume additivity, and thus may be more robust if the outcome model for (2) is misspecified. However, while OM, IPW1, and IPW2 are all consistent, OM is most efficient if correctly specified. Thus, depending on the availability of the data or the ability to estimate certain key quantities, a researcher may prefer one approach over the others.
@@
Regarding the incorporation of propensity and sample selection weights into the analysis, our weights are not the same as those in the traditional g-computation or IPTW frameworks. We derived a new estimator with weights both for within-sample inference and for generalizability to our target population. Specifically, note the term $\Pr(S=1\mid X)$ in the denominators of Formulas (2) and (4). If selection does not depend on $A$, then identification reduces to the standard proof for the IPTW identification formula. However, in our setting, selection depends on the participant’s self-reported race. Thus, we need the following decomposition to calculate $\Pr(S=1\mid X)$:
$\Pr(S=1\mid X)=\Pr(S=1\mid A=1,X)\Pr(A=1\mid X)+\Pr(S=1\mid A=0,X)\Pr(A=0\mid X).$
@@
As with OM, IPW2 requires that the probability of selection be marginalized over $A$, i.e., $\Pr(S=1\mid X)$, and that the propensity model be fit without the incorporation of selection weights, i.e., $\Pr(A=a\mid S=1,X)$. IPW1 requires the selection mechanism be known (e.g., in the case of survey weights), conditional on $A$, i.e., $\Pr(S=1\mid A=a,X)$, and that a survey-weighted propensity model be fit, i.e., $\Pr(A=a\mid X)$. The key difference here comes with respect to how the joint propensity and selection mechanisms are factorized.

### v2

@@ 3.2 Proposed Estimators
We wish to draw inference on the population from which our analytic sample was selected. For $a\in\{0,1\}$, we define
$\mu(a)=\mathbb{E}_{\mathcal{P}}\left[\mathbb{E}(Y\mid A=a,\boldsymbol{X})\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right],$
where $g_{a}(\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})$ and $w(\boldsymbol{X})$ is a known design selection factor. The target of inference, the ACD, is given by ${\rm ACD}=\mu(1)-\mu(0)$. We make the following assumptions:
@@
    Positivity: $e_{a}^{\mathcal{Q}}(\boldsymbol{X}):=\mathcal{Q}(A=a\mid\boldsymbol{X})>0$ $\mathcal{Q}$-a.s. $\forall a\in\mathcal{A}$
@@
    Selection Positivity: $\pi_{a}(\boldsymbol{X}):=\Pr(S=1\mid A=a,\boldsymbol{X})>0$ on the support of $A,\boldsymbol{X}$
@@
    Weak Selection Exchangeability: $\mathbb{E}_{\mathcal{P}}(Y\mid A=a,\boldsymbol{X})=\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})=g_{a}(\boldsymbol{X})$.
We require (i) and (ii) to ensure that our observed data functional, $\mu(a)$, is well-defined. Condition (iii) allows us to generalize the conditional mean of the outcome from the study sample to the target population. Specifically, the weak selection exchangeability assumption posits that the expected outcome, conditional on confounders, $\boldsymbol{X}$, and the exposure/group membership variable of interest, $A$, is the same between the sample and the target population. This assumption is less stringent than exchangeability in distribution (i.e, $Y\perp S\mid A,\boldsymbol{X}$), as it does not require the entire conditional distribution of the outcomes to be independent of selection, only their conditional expectations. Under survey sampling, this assumption automatically holds when the survey weights fully encode the (known) design and response adjustments. Further, when the parameter of interest is a conditional quantity, incorporation of the survey weights will not affect the bias or standard error if the model is correctly specified. However, we are interested in ACD, which is a marginal quantity, thus the weighting is needed to standardize to the target population. In other words, even though the survey weights adjust for the sampling design, incorporating them into the ACD is necessary to correctly weight the final estimator to generalize and deconfound estimation. How the survey weights are incorporated depends on the factorization of the joint probability of exposure/group membership and selection.
Depending on how we factorize the joint probability of being sampled $(S=1)$ and belonging to a particular group $(A=a)$, $\mu(a)$ admits the following equivalent identification formulas under the sample law, $\mathcal{Q}$, given by
@@
$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\mathbb{E}_{\mathcal{Q}}(Y\mid A=a,\boldsymbol{X})\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})g_{a}(\boldsymbol{X})\right],\quad\text{or}$ (1)
$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y\Pr(S=1)}{\Pr(S=1,A=a\mid\boldsymbol{X})}\right],$ (2)
where $w(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}=\tfrac{\Pr(S=1)}{\pi(\boldsymbol{X})}$, $\mathbb{I}(\cdot)$ is the indicator function, and Equation (2) is equivalent to the following two factored forms:
@@
$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y}{\mathcal{P}(A=a\mid\boldsymbol{X})}\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid A=a,\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w_{a}(\boldsymbol{X})\frac{\mathbb{I}(A=a)Y}{e_{a}^{\mathcal{P}}(\boldsymbol{X})}\right],\quad\text{or}$ (3)
$\mu(a)$ $=\mathbb{E}_{\mathcal{Q}}\left[\frac{\mathbb{I}(A=a)Y}{\mathcal{Q}(A=a\mid\boldsymbol{X})}\cdot\frac{\Pr(S=1)}{\Pr(S=1\mid\boldsymbol{X})}\right]=\mathbb{E}_{\mathcal{Q}}\left[w(\boldsymbol{X})\frac{\mathbb{I}(A=a)Y}{e_{a}^{\mathcal{Q}}(\boldsymbol{X})}\right],$ (4)
where $e_{a}^{\mathcal{Q}}(\boldsymbol{X}):=\mathcal{Q}(A=a\mid\boldsymbol{X})$ is the within-sample propensity, $e_{a}^{\mathcal{P}}(\boldsymbol{X}):=\mathcal{P}(A=a\mid\boldsymbol{X})$ is the population propensity, and $w_{a}(\boldsymbol{X})=\tfrac{\Pr(S=1)}{\Pr(S=1\mid A=a,\boldsymbol{X})}$. We refer to (1), (3), and (4) as Outcome Modeling (OM), Inverse Probability Weighting 1 (IPW1), and IPW2, respectively. For the full derivation of (1) - (4), see Appendix A.1.
@@
The assumptions required to estimate the ACD are relatively weaker than those required to identify a causal quantity. Figure 1, Panel (C) gives an example of a directed acyclic graph (DAG) where a causal target may be warranted, such as the case where $A$ is an environmental exposure, like lead, which can also alter telomere length. We provide a supplementary analysis in which we estimate the causal effect of lead exposure on telomere length for illustration, as a practitioner may also be interested in such causally interpretable estimated differences. If the group variable under study is a treatment or exposure, we show under stronger assumptions that Formulas (1) - (4) can equivalently target the population potential outcome means, denoted by $\mathbb{E}_{\mathcal{P}}[Y^{a}]\ \forall a\in\mathcal{A}$. Along with the previous assumptions (i) and (ii), the following conditions,
  1. (iii’)
    Weak Selection Exchangeability: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]=\mathbb{E}_{\mathcal{Q}}[Y^{a}\mid A=a,\boldsymbol{X}]$
@@
    Stable Unit Treatment Value Assumption (SUTVA): $Y_{i}=Y_{i}^{a}\ \forall i,\ A_{i}=a\in\mathcal{A}$
@@
    Weak Exchangeability: $\mathbb{E}_{\mathcal{P}}[Y^{a}\mid\boldsymbol{X}]=\mathbb{E}_{\mathcal{P}}[Y^{a}\mid A=a,\boldsymbol{X}]$,
are sufficient to identify $\mathbb{E}_{\mathcal{P}}[Y^{a}]$. SUTVA underpins the existence of the counterfactual outcomes, $Y^{a}$. The compact representation in (iv) captures two aspects of SUTVA – no interference, which states that the $i$th observation’s outcome is unaffected by the treatment of other observations and consistency, which states that the potential outcome is well-defined for each treatment level, $a$ (Rubin, 1980). The weak exchangeability assumption allows us to represent the conditional distribution of the unobserved potential outcome using that of the observed potential outcome. As before, we only require a weaker version of this assumption, though strong exchangeability, i.e., $Y^{a}\perp A\mid\boldsymbol{X}$, is often assumed. For additional details and the full identification proof, see Appendix A.2. This is thematically similar to Dahabreh et al. (2020), who study the generalization of causal inferences from individuals in randomized trials. The authors also differentiate between mean exchangeability in $A$ for the randomized population and mean exchangeability in $S$ for the target population in the context of randomized trials.
@@ 3.3 Comparisons to Existing Approaches
Formulas (1) - (4) differ from standard approaches one might take in practice, which warrants a brief discussion here. Table 1 compares our proposed approaches to such existing methods. If our assumed conceptual diagram holds, there are two sources of bias: bias from covariate imbalance between comparison groups (i.e., confounding bias) and unequal selection induced by the design of the complex survey (i.e., selection bias), which impacts the generalizability of our within-sample results to a larger, target population. Existing approaches may use balancing and/or generalizability weights to address these biases, but omitting either can result in biased estimates of our target quantity. For instance, the effect of race can be either exaggerated or attenuated depending on the direction of the true association between confounders, race, and the outcome in simple regression (Approach 1 in Table 1; Williams et al., 2019). Multiple regression (Approach 2) adjusts for covariates but will suffer from bias due to model misspecification unless the outcome model includes, for example, appropriate interaction terms (Rao, 1971). With inverse probability weighting (Approach 3), the race-covariate relationship is modeled, rather than the covariate-outcome relationship, foregoing the need to specify any race $\times$ covariate interactions (Mansournia and Altman, 2016). However, with all three of these approaches, the target of inference is the within-sample difference in outcomes, which cannot be generalized to a larger population if a probability sample was taken (Shi et al., 2023; Degtiar and Rose, 2023). Survey-weighted regression (Approach 4) accounts for selection bias by accounting for how many individuals in the population that one individual in the sample represents (Lumley and Scott, 2017), allowing for generalizability, but will be biased if there is covariate imbalance (DuGoff et al., 2014).
@@
Table 1: Comparison of existing approaches for analysis to our proposed approach. Dashes (–) denote that the model is not used.
Model Bias Being Addressed
Approach Outcome Propensity Selection Confounding Selection
Existing Approaches
1. Simple Regression $\mathbb{E}[Y\mid A,S=1]$ No No
2. Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ No No
3. IPTW Estimatora $\Pr(A=1\mid\boldsymbol{X},S=1)$ Yes No
4. Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(S=1\mid A,\boldsymbol{X})$ No Yes
5. IPTW Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ Yes No
6. IPTW + Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
7. Weighted IPTW + Survey-Weighted Multiple Regression $\mathbb{E}[Y\mid A,\boldsymbol{X}]$ $\Pr(A=1\mid\boldsymbol{X})$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
Proposed Approaches
8. Outcome Modeling and Direct Standardization $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(S=1\mid\boldsymbol{X})^{{\bf b}}$ Yes Yes
9. Inverse Probability Weighting 1 $\Pr(A=1\mid\boldsymbol{X})$ $\Pr(S=1\mid A,\boldsymbol{X})$ Yes Yes
10. Inverse Probability Weighting 2 $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid\boldsymbol{X})$ Yes Yes
11. Augmented Inverse Probability Weighting $\mathbb{E}[Y\mid A,\boldsymbol{X},S=1]$ $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid\boldsymbol{X})$ Yes Yes
a IPTW: Inverse Probability of Treatment Weighting.
b $\Pr(S=1\mid\boldsymbol{X})=\Pr(S=1\mid A=1,\boldsymbol{X})\Pr(A=1\mid\boldsymbol{X})+\Pr(S=1\mid A=0,\boldsymbol{X})\Pr(A=0\mid\boldsymbol{X})$.
@@ 10. Inverse Probability Weighting 2 $\Pr(A=1\mid\boldsymbol{X},S=1)$ $\Pr(S=1\mid\boldsymbol{X})$ Yes Yes
In our research question, we have an intersecting consideration of both selection and confounding bias; thus, neither source of bias should be disregarded. A more principled approach would be one that accounts for both confounding and selection bias, as given by Approaches 6 and 7, which utilize a combination of propensity score and survey weighting. These approaches, advocated by Zanutto (2006), DuGoff et al. (2014), and Ridgeway et al. (2015) are considered the current analytic standard. They suggest survey-weighted generalizations of IPTW, either at the propensity model stage and/or the outcome model stage. Although it may seem that either extension addresses both sources of bias, there is little theoretical justification for how weighting one or multiple models within IPTW achieves both tasks in our setting. Additionally, propensity scores are often seen as in-sample balancing scores (Zanutto, 2006; DuGoff et al., 2014). From this perspective, their generalizability (and corresponding selection bias correction based on survey weights) is often ignored. Survey-weighting both the propensity score and outcome model is reasonable (Ridgeway et al., 2015), however directly combining survey-weighted IPWs and survey weights in a second stage outcome regression may also lead to biased results.
@@
From a practical standpoint, one question is: ‘should we implement survey-weighted propensity score estimation in IPTW or not?’ Our analytic results show that for estimation of the ACD, the propensity score does not have to be survey-weighted. We see this in Equations (3) and (4), which differ with respect to the factorization of the joint distribution into propensity and survey weight models. If the propensity score is not survey-weighted, based on Equation (4), we need $\Pr(S=1\mid\boldsymbol{X})$, i.e., one may need to marginalize $\Pr(S=1\mid A,\boldsymbol{X})$ over the group membership variable, $A$. If survey weights are included in the propensity score estimation, then we must know $\Pr(S=1\mid A,\boldsymbol{X})$, i.e., the probability of selection conditional on each level of the group membership variable and the covariates.
@@
Moreover, our proposed approaches differ from each other and prove useful in different settings. Formula (1), outcome modeling and direct standardization (OM), requires knowledge of $f_{X}(x)$, as an outcome regression, $g_{a}(\boldsymbol{X})$, is fit in practice. Formulas (3) and (4), IPW1 and IPW2, do not require estimation of an outcome model and may be the best options depending on the level of information available in practice. We view IPW1 and IPW2 as simple corrections to what one might do in practice, namely the traditional inverse probability of treatment weighted (IPTW) estimator, and thus may be a better starting point for an analysis of this type as they additionally account for selection bias. While IPW1 does align better with the NHANES survey weighting structure, IPW1 versus IPW2 show that the incorporation of the survey weights into each stage of modeling depends on two alternate factorizations for the joint distribution of being sampled and belonging to the exposure/group of interest. The choice of IPW1 versus IPW2 is also driven by design, as IPW1 is better suited for retrospective designs, where IPW2 is better suited for prospective designs. Further, IPW1 and IPW2 may be more robust if the outcome model for (1) is likely misspecified, however, the additional modeling of the two propensity score models for IPW2 may result in efficiency loss. While OM, IPW1, and IPW2 are all consistent, OM is most efficient if correctly specified. Thus, depending on the availability of the data or the ability to estimate certain key quantities, a researcher may prefer one approach over the others. A doubly robust estimator, by combining both an outcome model and an inverse-probability weighting component, requires only one of these two models to be correctly specified in order to yield a consistent ACD estimate (see Section 3.4.1). In practice, this flexibility eliminates the need for a choice strictly between IPW1 and IPW2. If either the propensity model or the outcome model are well-specified, the doubly robust approach remains consistent and often achieves smaller variance than either IPTW estimator alone. For these reasons, we generally recommend the doubly robust estimator as a practical default.
@@
Regarding the incorporation of propensity and sample selection weights, our weights differ from those in traditional g-computation or IPTW frameworks. We derived a new estimator with weights for both within-sample inference and generalizability to our target population. Specifically, note the term $\Pr(S=1\mid\boldsymbol{X})$ in the denominators of Formulas (1) and (4). If selection does not depend on $A$, identification simplifies to the standard IPTW identification formula. However, in our setting, selection depends on the participant’s self-reported race, requiring the probability of selection to be marginalized over $A$, where
$\Pr(S=1\mid\boldsymbol{X})=\Pr(S=1\mid A=1,\boldsymbol{X})\Pr(A=1\mid\boldsymbol{X})+\Pr(S=1\mid A=0,\boldsymbol{X})\Pr(A=0\mid\boldsymbol{X}).$
@@
As with OM, IPW2 uses the marginalized $\Pr(S=1\mid\boldsymbol{X})$ above, rather than directly modeling $\Pr(S=1\mid\boldsymbol{X})$ by omitting $A$ (Wilms et al., 2021). Thus, IPW2 requires that both the propensity model be fit without incorporating selection weights, i.e., $e_{a}^{\mathcal{Q}}(\boldsymbol{X})$, for inverse probability weighting, and the survey-weighted propensity model be fit to calculate $\Pr(S=1\mid\boldsymbol{X})$. In contrast, IPW1 requires the selection mechanism be known (e.g., in the case of survey weights), conditional on $A$, i.e., $\Pr(S=1\mid A=a,\boldsymbol{X})$, and that a survey-weighted propensity model be fit, i.e., $e_{a}^{\mathcal{P}}(\boldsymbol{X})$. The key difference here comes with respect to how the joint propensity and selection mechanisms are factorized.